# footnotetext: footnotemark: Work done as Student Researcher at Google DeepMindfootnotetext: footnotemark: Work completed prior to joining Apple

### v1

@@ Appendix A Feature Extraction
To comprehensively analyze the complexity of features extracted from a deep learning model, we employed a detailed feature extraction process using dictionary learning, specifically utilizing an over-complete dictionary. This approach allows each activation $\bm{f}_{n}(\mathstatbold{x})\in\mathcal{A}_{\ell}$ to be expressed as a linear combination of multiple basis elements $\bm{v}\in\mathcal{A}_{\ell}$ (atoms) from the dictionary $\mathstatbold{D}^{\star}=\{\bm{v}_{1},\ldots,\bm{v}_{k}\}$ coupled with some sparse coefficient $\bm{z}\in\mathbb{R}^{k}$ associated to each atoms.
@@
$\bm{f}_{n}(\mathstatbold{x})\approx\bm{z}\mathstatbold{D}^{\star}=\sum_{i}^{k}z_{i}\bm{v}_{i}$
where $z_{i}$ are the coefficients indicating the contribution of each atom $\bm{v}_{i}$ from the dictionary.
@@ Appendix C Complexity measure
Given that the input data are centered and scaled, we typically have $\text{Var}(\mathstatbold{x})=1$ at the input layer. Furthermore, residual connections and batch normalization tend to preserve this scaling in deeper layers, implying $\text{Var}(\mathstat{z})\approx 1$.
We define complexity as the opposite of the average $\mathcal{V}$-information across layers: complex features are those that are harder to decode. We add a shift of $1$ for the ease of plotting. Empirically, we observed that this adjustment yields $K(\mathstat{z},\mathstatbold{x})$ in the range $[0,1]$, with $1$ indicating a complex feature that is not available and 00 indicating a simple feature that is fully available.
@@ Appendix G Importance Measure
In our case, we measured the importance by taking the absolute value of the importance vector, i.e., $\bm{\Gamma}(\mathstat{z}_{i})=\mathbb{E}_{\mathbb{P}{z}}\Big{(}||\frac{\partial\mathstat{y}}{\partial\mathstat{z}_{i}}\cdot\mathstat{z}_{i}||\Big{)}$. It is natural to question whether this approach might overlook important features due to their inhibitory effects. Indeed, as depicted in Figure 12, a large number of features may be important not because they add positive energy to the logits, but by inhibition, i.e., by suppressing class information. Although this does not alter the implications of our previous observations, it is noteworthy that the majority of inhibitory features are also simple features.
@@
Another property of importance is its close relationship with prevalence [30], which indicates that a frequently occurring feature will, on average, be more important given the same importance coefficient ($\nabla_{\mathstat{z}}\mathstat{y}$). In our study, this implies that if the most important features are reduced, these important features are also potentially more frequently present. Consequently, the prevalence of a feature can be a factor explaining this sedimentation process. We refer the reader to a concurrent study that proposed to investigate more deeply this phenomena using controlled dataset [55].
Appendix H Kolmogorov, Levin and $\mathcal{V}$-information

### v2

# footnotetext: **footnotemark: * These authors contributed equally to this workfootnotetext: footnotemark: Work done as Student Researcher at Google DeepMindfootnotetext: footnotemark: Work completed prior to joining Apple

@@ Appendix A Feature Extraction
To comprehensively analyze the complexity of features extracted from a deep learning model, we employed a detailed feature extraction process using dictionary learning, specifically utilizing an over-complete dictionary. This approach allows each activation $\bm{f}_{n}(\mathstatbold{x})\in\mathcal{A}_{\ell}$ to be expressed as a linear combination of multiple basis elements (direction, also called atoms) $\bm{d}\in\mathcal{A}_{\ell}$ from the dictionary $\mathstatbold{D}^{\star}=\{\bm{d}_{1},\ldots,\bm{d}_{k}\}$ coupled with some sparse coefficient $\bm{z}\in\mathbb{R}^{k}$ associated to each atoms.
@@
$\bm{f}_{n}(\mathstatbold{x})\approx\bm{z}\mathstatbold{D}^{\star}=\sum_{i}^{k}z_{i}\bm{d}_{i}$
where $z_{i}$ are the coefficients indicating the contribution of each atom $\bm{d}_{i}$ from the dictionary.
@@ Appendix C Complexity measure
Given that the input data are centered and scaled, we typically have $\text{Var}(\mathstat{z})$ around 1 (or less in case of Layer normalization). Furthermore, residual connections and batch normalization tend to preserve this scaling in deeper layers. We note, however, that our score $K(\mathstat{z},\mathstatbold{x})$ is not strictly bounded. Indeed, we define complexity as the opposite of the average $\mathcal{V}$-information across layers: complex features are those that are harder to decode. We add a shift of $1$ for the ease of plotting. Empirically, we observed that this adjustment yields $K(\mathstat{z},\mathstatbold{x})$ in the range $[0,1]$, with $1$ indicating a complex feature that is not available and 00 indicating a simple feature that is fully available.
@@ Appendix G Importance Measure
In our case, we measured the importance by taking the absolute value of the importance vector, i.e., $\bm{\Gamma}(\mathstat{z}_{i})=\mathbb{E}_{\mathbb{P}{z}}\Big{(}||\frac{\partial\mathstat{y}}{\partial\mathstat{z}_{i}}\cdot\mathstat{z}_{i}||\Big{)}$. It is natural to question whether this approach might overlook important features due to their inhibitory effects. Indeed, as depicted in Figure 12, numerous features may be important not because they add positive energy to the logits, but by inhibition, i.e., by suppressing class information. Although this does not alter the implications of our previous observations, it is noteworthy that the majority of inhibitory features are also simple features.
@@
Another property of importance is its close relationship with prevalence [32], which indicates that a frequently occurring feature will, on average, be more important given the same importance coefficient ($\nabla_{\mathstat{z}}\mathstat{y}$). In our study, this implies that if the most important features are reduced, these important features are also potentially more frequently present. Consequently, the prevalence of a feature can be a factor explaining this sedimentation process. We refer the reader to a concurrent study that proposed to investigate more deeply this phenomena using a controlled dataset [58].
Appendix H Features Clustering
The visualization in Figure 2 prompts an important question: are features in our model clustered based on their complexity? Specifically, are there regions in the feature space that are generally more “complex” than others, or that respond primarily to more complex stimuli? Our access to 10,000 features via overcomplete decomposition enables a more detailed analysis compared to traditional neuron-wise studies. We aim to explore three main hypotheses:
  • Hypothesis 1: Features cluster by super-class. This hypothesis posits that features corresponding to semantically related categories are spatially grouped within the feature space—e.g., concepts related to the dog class are closer to those of the cat class than to unrelated classes like furniture.
  • Hypothesis 2: Features cluster by complexity. We suggest that features may organize themselves based on their complexity, with simpler features forming distinct clusters separate from more complex ones.
  • Hypothesis 3: Features cluster by importance. This hypothesis explores whether features with similar predictive importance tend to group together within the feature space.
It is important to emphasize that these hypotheses are mutually independent. To test them, we propose two methodologies: (1) visualizing feature embeddings using UMAP and (2) clustering the features followed by dendrogram analysis to examine whether the resulting clusters are homogeneous (also called "pure") in terms of super-class, complexity, or importance.
Figure 13: Feature Similarity vs Super-Class. Each point represents a concept, with its color indicating the associated super-class. Some super-classes such as birds, reptiles, dogs & other mammals form well-defined, tight clusters, suggesting that features belonging to them are close in the feature space. Others, such as device, clothes appear more dispersed. By comparing this figure with Figure 2, we can identify which meta-features are "pure" (belonging to a single super-class) and which are "impure" (spanning multiple super-classes). Interestingly, the “impurity” region seems to cover low-complexity and mid-complexity concepts such that Grass, Waves, Trees, Low-pixel quality detector which are not class-specific.
Feature Similarity vs Super-Class.
Figure 13 illustrates the organization of features by their super-class. Each point is a feature, colored according to its super-class label. We observe that certain super-classes, like those associated with birds, dogs, reptiles form distinct and cohesive clusters, indicating a strong grouping within the feature space. Other region have features that encompassing a wider range of super-class, such that grass, waves, low-pixel quality detector. This reveals that some meta-features are predominantly associated with a single super-class ("pure"), while others span multiple super-classes ("impure"), reflecting the shared visual characteristics or multi-functional nature of those features.
Figure 14: Feature Similarity by Complexity. A) Each point is a feature, colored by its complexity score. Distinct areas of the graph correspond to varying levels of complexity, suggesting a non-random distribution of feature complexity. For instance, animal-related features tend to have higher complexity, one could hypothesize that the fine-grained classification required for these categories are responsible for this complexity. B) A four-level dendrogram where each level further segments clusters and calculates the average complexity for each sub-cluster. A clear split by complexity appears at the first level and intensifies with depth, supporting the idea that some regions of the feature space are inherently more complex than others.
Clustering by Complexity.
Figure 14 explores how features are organized based on their complexity. Panel A shows a UMAP visualization with points colored according to their complexity scores. The visualization reveals distinct regions of varying complexity, indicating that the distribution is structured. For instance, features related to animals display higher complexity, likely because these require fine-grained and precise detectors. In contrast, simpler features, such as color detectors or low-frequency patterns, cluster together in less complex regions. Panel B displays a dendrogram with four hierarchical levels, where each level introduces additional splits, and the mean complexity is calculated for each sub-cluster. The pronounced division in complexity at the first level, which sharpens as we delve deeper, suggests that the feature space is compartmentalized based on complexity. A promising direction for future research would be to align these complexity clusters with known visual cortical areas to explore potential correspondences.
Figure 15: Feature Similarity by Importance. A) Each point represents a feature, with color indicating its importance. Distinct regions of the graph contain features of varying importance, particularly with more important features clustering at the top. B) A four-level dendrogram with sub-clusters evaluated by their average importance. We observe that from the first level, the dendrogram effectively splits features into groups of varying importance, corresponding to the upper and lower parts of the UMAP graph in Panel A.
Clustering by Importance.
Finally, we investigate the third hypothesis: the presence of regions within the feature space that contain features with higher predictive importance. Figure 15 Panel A depicts a UMAP visualization where features are colored by their importance. The graph shows a clear structure, with highly important features grouping in specific regions (e.g., the upper part of the graph), while less important features are distributed in other regions. Panel B provides the corresponding dendrogram, revealing that even at the first level of the hierarchy, features segregate into clusters of varying importance. Notably, low-complexity “support” features—such as grass, waves, and low-pixel-quality detectors—tend to form cohesive clusters. Meanwhile, more predictive features, like animal-related features that drive classification, group together in another distinct region. This raises a crucial question: is this clustering merely correlated (i.e., based on shared visual aspects like context or background), or does it reflect a causal relationship in the model’s predictive structure? This question remains an open avenue for future investigation.
Appendix I Local vs Distributed Encoding
Figure 16: Local vs Distributed Encoding. Each point represents a feature, with color indicating its Hoyer score. Higher scores suggest a more "local" representation, where a feature is primarily encoded by a single neuron. Lower scores indicate a distributed representation across a population of neurons. Interestingly, some features have scores near 1, implying near-complete localization, while others are more distributed. This variation highlights the diversity in encoding across features.
Neural networks exhibit a diversity in how features are encoded, ranging from local to distributed representations [25, 29]. In the local encoding scenario, a single neuron is primarily responsible for encoding a feature. On the other hand, in a distributed encoding scheme, features are represented by the coordinated activity of multiple neurons, often deviating from canonical axis-aligned directions. More specifically, features could be distributed across many neurons—sometimes densely, or in a pseudo-distributed fashion—rather than being localized to a specific neuron. This theory is one of the main motivations behind employing overcomplete concept extraction methods [18, 32, 92, 35].
The distinction between local and distributed encoding is critical, particularly when extracting overcomplete features from a large-scale dictionary. In our analysis of 10,000 features, we assess whether these features are encoded in a local or distributed manner. A local encoding would imply that the feature vector $\bm{d}$ is aligned with a canonical vector, i.e., $\bm{d}\in\{\bm{e}_{1},...,\bm{e}_{n}\}$, where $n=|\mathcal{A}_{\ell}|$ is the dimensionality of the activation space and $\bm{e}_{i}$ represents the one-hot canonical vector for the $i$th neuron. In contrast, a fully distributed feature would be characterized by non-zero values across all neurons, with no alignment to any single axis.
To quantify the extent of this local or distributed encoding, we use the Hoyer score. This score, which ranges between 0 and 1, captures the sparsity of a vector by comparing its $\ell_{1}$-norm to its $\ell_{2}$-norm, with a correction term for normalization. Formally, the Hoyer score is given by:
$\text{Hoyer}(\bm{d})=\frac{\sqrt{n}-||\bm{d}||_{1}/||\bm{d}||_{2}}{\sqrt{n}-1}.$
For each feature in our dictionary $\mathstatbold{D}=\{\bm{d}_{1},...,~{}\bm{d}_{10,000}\}$, we compute the Hoyer score to determine whether the feature is locally encoded (score near 1) or distributed (score near 0).
Figure 16 illustrates the results of this analysis, revealing significant variability in the degree of distribution among features. Some regions of the feature space show highly distributed encoding, while others exhibit more local representations. This analysis was conducted on a ResNet50 model, and it is possible that the distribution of encoding strategies varies across different architectures.
Figure 17: Feature Complexity vs. Distributed Encoding. We show that there is no clear relationship between feature complexity and the degree of distributed encoding. Whether a feature is encoded by a single neuron or distributed across multiple neurons does not seems to be determined by its complexity.
Appendix J Replicating the feature flow results with other measures and models.
Figure 18: Replication of Figure 4 Using $\mathcal{V}$-Information. As described in Section 4, we replicate the analysis of feature flow and complexity using $\mathcal{V}$-information as a measure.
A legitimate consideration arises when revisiting Section 4, where we introduced the hypothesis that simpler features are primarily carried through the residual connections of a network, while more complex features are progressively constructed through interactions between the main branch and residual connections. The measure we originally used to support this hypothesis was CKA, which serves as a proxy to assess the similarity between the activations at a certain stage of the model and the final state of a concept. However, one might wonder why not use $\mathcal{V}$-information directly.
Figure 18 presents the results of this replication. Although the scale of the values differs from those in Section 4, the overall trend remains consistent. The left panel shows the dynamics of simpler features. Early in the network, the main branch carries a significant portion of the $\mathcal{V}$-information, which diminishes as the simpler features are gradually “passed along” through the residual connections. Notably, even as the information content decreases, the absolute quantity of $\mathcal{V}$-information remains higher for simple features compared to complex ones. This indicates that while simpler features are transported through the residual connections, they are not entirely depleted of their information content. The right panel of Figure 18 demonstrates the progressive construction of more complex features. Here, we see that both the main branch and residual connections contribute to the gradual accumulation of information necessary for these complex features. This supports our initial hypothesis: complex features do not traverse the network intact but are built up in a cumulative process, drawing on multiple layers and branches to form intricate representations.
For compute consideration, this replication was conducted using a different experimental setup than that of Section 4. For this analysis, we employed the validation set of ImageNet to build the dictionary instead of the train set. Both the dictionary and the model were different from those used in the main body of the paper. Specifically, the model used here was the ResNet50 implementation from the Keras library [26]. Despite these differences in experimental conditions, the overarching trends observed in our original CKA-based analysis are preserved, bolstering the validity of the flow hypothesis.
Appendix K Kolmogorov, Levin and $\mathcal{V}$-information